OpenStudy (anonymous):

-37 + (-5) (1 point)

OpenStudy (anonymous):

-42

OpenStudy (saifoo.khan):

-42

OpenStudy (anonymous):

-42

OpenStudy (anonymous):

-37+(-5) = -37-5 = -42

OpenStudy (anonymous):

-42